OpenStudy (anonymous):

Find the quotient. (35n3 - 30n2 + 25n) ÷ (-5n) a) -7n^2 - 6n + 5 b) -7n^2 - 6n - 5 c) -7n^2 + 6n - 5

OpenStudy (anonymous):

The anwer is c. You would write the left hand expression (35n3-30n2+25n) in the numerator and the right hand expression (-5n) in the denominator. You can then separate the terms because they are added by taking each term in the numerator and putting it over the term in the denominator. \[(35n ^{3})/(-5n) = -7n^{2}\] \[(-30n^{2})/(-5n) = 6n\] \[(25n)/(-5n) = -5 \] Add each term and that is the answer.